Consider the following recurrence T(n) = 3T(n/5) + c What is the complexity of T(n) in...

  1. Consider the following recurrence
  2. T(n) = 3T(n/5) + c

    What is the complexity of T(n) in Big O? ___________.
Answer #1

Time Complexity

g T(n)= 37(415) + c .-. for Put n = n15 T(n)) = 3 T (hlas) + c T(n)= 3 (3 T(2125)+) + C. = 34T (W125) +Be+ @ put h=h125 iner=3h T (1154) + 3k 1 ( +34-2 ( to t34 +3cte kth team for C higk=1 n=5k log both sider, logn= klogs lo5х = к lo95 K=logh 3°C+3